# Searching for Inflationary Physics with the CMB Trispectrum:

### v2

@@ I Introduction
This work presents the PolySpec code: a fast and flexible package for performing quasi-optimal analyses of CMB temperature and polarization data. This heavily extends the PolyBin code presented in [44, 45, 54] (which estimates binned power spectra, bispectra and trispectra), and allows for minimum-variance estimation of eleven types of primordial trispectrum as well as two late-time effects. In particular, we build estimators for quadratic and cubic local non-Gaussianity (parametrized by $g_{\rm NL}^{\rm loc}$ and $\tau_{\rm NL}^{\rm loc}$), constant non-Gaussianity ($g_{\rm NL}^{\rm con}$), the three self-interaction EFTI templates ($g_{\rm NL}^{\dot{\sigma}^{4}},g_{\rm NL}^{\dot{\sigma}^{2}(\partial\sigma)^{2}},g_{\rm NL}^{(\partial\sigma)^{4}}$), direction-dependent exchange non-Gaussianity assuming a parity-even ($\tau_{\rm NL}^{n,\rm even}$), parity-odd ($\tau_{\rm NL}^{n,\rm odd}$) or generalized ($\tau_{\rm NL}^{n_{1}n_{3}n}$) basis set, intermediate annd light spin-$s$ field exchange ($\tau_{\rm NL}^{\rm light}(s,\nu_{s})$), heavy spin-$s$ field exchange ($\tau_{\rm NL}^{\rm heavy}(s,\mu_{s})$), gravitational lensing $(A_{\rm lens}$) and unresolved point sources ($t_{\rm ps}$). [^1] Precise template definitions can be found in Paper 1 and are summarized in the appendix of Paper 3; here, we note the rough form and features of each template when it is introduced in §IV. The PolySpec code is built in python for flexibility and makes extensive use of the ducc package for spherical harmonic transforms (which is an upgrade to libsharp [55]), with additional rate-limiting steps implemented in c. In addition to presenting the code, we perform an extensive battery of validation tests, affording us confidence that (a) the estimators are working as expected, (b) the various hyperparameters are well chosen, and (c) the code makes efficient use of computational resources.
The remainder of this work is as follows. In §II, we give an overview of the trispectrum estimators introduced in Paper 1, paying close to attention to their statistical properties and underlying approximations. Next, §III presents the PolySpec code, including details of our numerical implementation and its basic usage. An extensive suite of forecasts and validation tests are performed in §IV, assessing the bias and optimality of each class of estimator under various conditions, in addition to the dependence on a variety of hyperparameters. In §V, we assess the computational scalings and runtimes of PolySpec, before concluding in §VI with a summary and a plan for the final installment of this series: application to the Planck dataset. For reference, we summarize the hyperparameters used in the estimators in Tab. 1 and demonstrate the basic usage of PolySpec in Fig. 1.
@@ II.1 Definition & Properties
We begin with a brief summary of the trispectrum estimators discussed in Paper 1, paying close attention to the various numerical approximations that will be tested in this work. Our starting point is the four-point function of the primordial curvature perturbation, $\zeta$, expressed in terms of a set of amplitudes $\{A_{\alpha}\}$ and corresponding templates $T^{(\alpha)}_{\zeta}$:
@@ III The PolySpec Code
k_array, transfer_array, lmin=lmin, lmax=lmax, Lmin=Lmin, Lmax=Lmax)
@@ IV.3.3 Gaussian Bias
Figure 3: Breakdown of contributions to the $g_{\rm NL}^{\rm loc}$ (left) and $\tau_{\rm NL}^{\rm loc}$ (right) trispectrum estimators applied to Gaussian (left sub-panels) and non-Gaussian (right sub-panels) simulations, with fiducial values indicated by the titles. We show the quartic (red), quadratic (blue), and constant (green) numerator contributions normalized by the inverse Fisher matrix, as well as their sum (black), which gives the full estimator. As expected, the three terms precisely cancel if the fields are Gaussian, but reveal subtle signals consistent with the injected values (shown in dotted lines) in the non-Gaussian case. Results are generated from the mean of 50 ($g_{\rm NL}^{\rm loc}$) and $100$ ($\tau_{\rm NL}^{\rm loc}$) masked temperature and polarization simulations for four choices of $\ell_{\rm max}$ and errorbars showing the error in a single realization. We perform a detailed analysis of the injected signals in Fig. 7.
@@ IV.3.7 Constant Non-Gaussianity
Figure 9: Contributions to the ideal $g_{\rm NL}^{\dot{\sigma}^{4}}$ Fisher matrix diagonal as a function of the integration variables $r,\tau$ (generalizing Fig. 9). The integrand is sharply peaked around $r=r_{\star}$ with $-\tau\lesssim r_{\star}$, and its full structure can be approximated using $35$ points, shown in red dots. The horizontal stripe around $r\approx 0.7r_{\star}$ corresponds to reionization. This is computed for using $T$- and $E$-modes for $\ell\in[2,512]$. Results for the other EFTI templates, $g_{\rm NL}^{\dot{\sigma}^{2}(\partial\sigma)^{2}}$ and $g_{\rm NL}^{(\partial\sigma)^{4}}$, are analogous.
@@ IV.6.2 Spin-Zero Results
We now validate the spin-zero estimators using Gaussian simulations. Our key results are shown in Fig. 17, comparing the empirical and theoretical errorbars on $\tau_{\rm NL}^{\rm light,heavy}$ across a wide range of mass parameters $\nu_{0}$ and $\mu_{0}$. The two sets of errorbars agree within $1.8\sigma$ (accounting for the expected scatter), validating our estimators. The addition of polarization leads to significantly tighter constraints, with $\sigma(\tau_{\rm NL})$ reducing by $20\%$ for local-like templates ($\nu\gtrsim 3/4$) and $40\%$ for the remaining templates.
@@
Figure 19: Dependence of the spin-zero collider constraints on the maximum internal mode, $L_{\rm max}$. We plot the ratio of the empirical (points) and theoretical (lines) errors as a function of $L_{\rm max}$ for fixed $\ell_{\rm max}=512$, using the same dataset as in Fig. 18 (without accounting for sample variance cancellation). In the conformal limit of $\nu_{0},\mu_{0}\to 0$, restricting to $L_{\rm max}<128$ causes significant loss of information, but we find negligible changes at larger $L_{\rm max}$. This implies that our $k$-space truncation restricts the analysis to collapsed configurations, as desired.
@@ IV.7 CMB Lensing
Finally, PolySpec includes estimators for two late-time effects: point-sources and gravitational lensing. Although they do not encode inflationary physics, these effects are important contaminants to primordial non-Gaussianity searches and can be accounted for using joint analyses [e.g., 11, 2]. The point-source effect is sourced by an uncorrelated Poisson distribution of sources, which produces an angle-independent trispectrum with amplitude $t_{\rm ps}$. As shown in Paper 1, the corresponding estimator is trivial (simply a quartic product of the inverse-variance-filtered data) and will discussed further in this work. The PolySpec lensing estimator takes a similar form to the $\tau_{\rm NL}^{\rm loc}$ estimator (involving the ‘exchange’ of a lensing field), and optimally estimates the lensing amplitude $A_{\rm lens}$ ($\propto C_{L}^{\phi\phi}$), using $T$-, $E$-, and $B$-modes. As described in Paper 1, this is analogous to the conventional quadratic lensing estimators [e.g., 85, 86, 87, 57], but naturally incorporates the $\mathrm{N}^{(1)}$ bias, first-principles ‘realization-dependent-noise’ [59], optimal polarization correlations [cf. 88], and mask- and weighting-dependent normalization.
@@
To validate the lensing estimator, we employ $150$ Gaussian realizations of temperature, polarization and lensing potential. [^19] Specifically, we use the FFP10 realizations [e.g., 89], which are publicly available on nersc. Using the lenspyx code [90], we generate non-Gaussian simulations with a known lensing amplitude $A_{\rm lens}$, which are supplemented with Gaussian noise and a Planck-mask, as before. Here, we use the hyperparameter set $\{N_{\rm disc}=100,N_{\rm fish}=50,\ell_{\rm min}=2,L_{\rm min}=2,N_{\rm side}=512\}$, fixing $L_{\rm max}=\ell_{\rm max}$ and using lensed $TT,TE,EE$ and $BB$ spectra where necessary [91], noting that the lensing estimators do not require optimization (since they do not contain radial integrals). Given that the fiducial lensing amplitude is non-zero, it is important to validate PolySpec at high resolution, thus we perform mock analyses up to $\ell_{\rm max}=2048$ (increasing $N_{\rm side}$ to $1024$ where necessary). [^20] As discussed above, it is safe to validate the other estimators at a reduced $\ell_{\rm max}$ since the errors scale with $\sigma(A)$ in the absence of a signal (assuming that the disconnected contributions are correctly subtracted).
@@ VI Summary & Next Steps
We close by elucidating our next steps. Up to this point, we have presented the inflationary templates and their optimal estimators (Paper 1) and built an efficient code to search for such signatures (this paper). In Paper 3, we will search for non-Gaussian signatures using Planck data. Motivated by the above results, we will perform both single-template and joint analyses of a range of inflationary templates, including all of the shapes considered in this work except for the poorly constrained direction-dependent forms with odd $n_{1},n_{3}$ (cf. §IV.5). We will further account for CMB lensing contamination using our $A_{\rm lens}$ template, and perform both temperature-only and joint temperature-and-polarization analyses. The result will be strong constraints on a wide variety of inflationary phenomena; these can be further improved using future CMB or large-scale structure datasets [e.g., 92, 93, 82].
@@ Acknowledgements.
We thank Giovanni Cabass, William Coulton, Adriaan Duivenvoorden, Sam Goldstein, Colin Hill and Maresuke Shiraishi for insightful discussions. OHEP is a Junior Fellow of the Simons Society of Fellows, and thanks Lludwig van Beethoven for auditory assistance. OHEP would also like to thank the Center for Computational Astrophysics for their hospitality across the multiple years this set of papers took to write. The computations in this work were run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation.

### v3

@@ I Introduction
This work presents the PolySpec code: a fast and flexible package for performing quasi-optimal analyses of CMB temperature and polarization data. This heavily extends the PolyBin code presented in [44, 45, 54] (which estimates binned power spectra, bispectra and trispectra), and allows for minimum-variance estimation of eleven types of primordial trispectrum as well as two late-time effects. In particular, we build estimators for cubic and quadratic-squared local non-Gaussianity (parametrized by $g_{\rm NL}^{\rm loc}$ and $\tau_{\rm NL}^{\rm loc}$), constant non-Gaussianity ($g_{\rm NL}^{\rm con}$), the three self-interaction EFTI templates ($g_{\rm NL}^{\dot{\sigma}^{4}},g_{\rm NL}^{\dot{\sigma}^{2}(\partial\sigma)^{2}},g_{\rm NL}^{(\partial\sigma)^{4}}$), direction-dependent exchange non-Gaussianity assuming a parity-even ($\tau_{\rm NL}^{n,\rm even}$), parity-odd ($\tau_{\rm NL}^{n,\rm odd}$) or generalized ($\tau_{\rm NL}^{n_{1}n_{3}n}$) basis set, intermediate annd light spin-$s$ field exchange ($\tau_{\rm NL}^{\rm light}(s,\nu_{s})$), heavy spin-$s$ field exchange ($\tau_{\rm NL}^{\rm heavy}(s,\mu_{s})$), gravitational lensing $(A_{\rm lens}$) and unresolved point sources ($t_{\rm ps}$). [^1] Precise template definitions can be found in Paper 1 and are summarized in the appendix of Paper 3; here, we note the rough form and features of each template when it is introduced in §IV. The PolySpec code is built in python for flexibility and makes extensive use of the ducc package for spherical harmonic transforms (which is an upgrade to libsharp [55]), with additional rate-limiting steps implemented in c. In addition to presenting the code, we perform an extensive battery of validation tests, affording us confidence that (a) the estimators are working as expected, (b) the various hyperparameters are well chosen, and (c) the code makes efficient use of computational resources.
The remainder of this work is as follows. In §II, we give an overview of the trispectrum estimators introduced in Paper 1, paying close attention to their statistical properties and underlying approximations. Next, §III presents the PolySpec code, including details of our numerical implementation and its basic usage. An extensive suite of forecasts and validation tests are performed in §IV, assessing the bias and optimality of each class of estimator under various conditions, in addition to the dependence on a variety of hyperparameters. In §V, we assess the computational scalings and runtimes of PolySpec, before concluding in §VI with a summary and a plan for the final installment of this series: application to the Planck dataset. For reference, we summarize the hyperparameters used in the estimators in Tab. 1 and demonstrate the basic usage of PolySpec in Fig. 1.
@@ II.1 Definition & Properties
We begin with a brief summary of the trispectrum estimators discussed in Paper 1, highlighting the various numerical approximations that will be tested in this work. Our starting point is the four-point function of the primordial curvature perturbation, $\zeta$, expressed in terms of a set of amplitudes $\{A_{\alpha}\}$ and corresponding templates $T^{(\alpha)}_{\zeta}$:
@@ III The PolySpec Code
lmin, lmax, k_array, transfer_array, Lmin, Lmax)
@@ IV.3.3 Gaussian Bias
Figure 3: Breakdown of contributions to the $g_{\rm NL}^{\rm loc}$ (left) and $\tau_{\rm NL}^{\rm loc}$ (right) trispectrum estimators applied to Gaussian (left sub-panels) and non-Gaussian (right sub-panels) simulations, with fiducial values indicated by the titles. We show the quartic (blue), quadratic (red), and constant (green) numerator contributions normalized by the inverse Fisher matrix, as well as their sum (black), which gives the full estimator. As expected, the three terms precisely cancel if the fields are Gaussian, but reveal subtle signals consistent with the injected values (shown in dotted lines) in the non-Gaussian case. Results are generated from the mean of 50 ($g_{\rm NL}^{\rm loc}$) and $100$ ($\tau_{\rm NL}^{\rm loc}$) masked temperature and polarization simulations for four choices of $\ell_{\rm max}$ and errorbars showing the error in a single realization. We perform a detailed analysis of the injected signals in Fig. 7.
@@ IV.3.7 Constant Non-Gaussianity
Figure 9: Contributions to the ideal $g_{\rm NL}^{\dot{\sigma}^{4}}$ Fisher matrix diagonal as a function of the integration variables $r,\tau$ (generalizing the left panel of Fig. 2). The integrand is sharply peaked around $r=r_{\star}$ with $-\tau\lesssim r_{\star}$, and its full structure can be approximated using $35$ points, shown in red dots. The horizontal stripe around $r\approx 0.7r_{\star}$ corresponds to reionization. This is computed for using $T$- and $E$-modes for $\ell\in[2,512]$. Results for the other EFTI templates, $g_{\rm NL}^{\dot{\sigma}^{2}(\partial\sigma)^{2}}$ and $g_{\rm NL}^{(\partial\sigma)^{4}}$, are analogous.
@@ IV.6.2 Spin-Zero Results
We now validate the spin-zero estimators using Gaussian simulations. Our key results are shown in Fig. 17, comparing the empirical and theoretical errorbars on $\tau_{\rm NL}^{\rm light,heavy}$ across a wide range of mass parameters $\nu_{0}$ and $\mu_{0}$. The two sets of errorbars agree within $1.8\sigma$ (accounting for the expected scatter), validating our estimators. The addition of polarization leads to significantly tighter constraints, with $\sigma(\tau_{\rm NL})$ reducing by $20\%$ for local-like templates ($\nu_{0}\gtrsim 3/4$) and $40\%$ for the remaining templates.
@@
Figure 19: Dependence of the spin-zero collider constraints on the maximum internal mode, $L_{\rm max}$. We plot the ratio of the empirical (points) and theoretical (lines) errors as a function of $L_{\rm max}$ for fixed $\ell_{\rm max}=512$, using the same dataset as in Fig. 18 (without accounting for sample variance cancellation). In the conformal limit of $\nu_{0}\to 0$ for light particles, as well as for heavy particles at all $\mu_{0}$, restricting to $L_{\rm max}<128$ causes significant loss of information, but we find negligible changes at larger $L_{\rm max}$. This implies that our $k$-space truncation restricts the analysis to collapsed configurations, as desired.
@@ IV.7 CMB Lensing
Finally, PolySpec includes estimators for two late-time effects: point-sources and gravitational lensing. [^19] PolySpec can additionally estimate trispectra arising from cross-correlations of CMB lensing and the integrated Sachs-Wolfe effect; this is discussed in detail in [85]. Although they do not encode inflationary physics, these effects are important contaminants to primordial non-Gaussianity searches and can be accounted for using joint analyses [e.g., 11, 2]. The point-source effect is sourced by an uncorrelated Poisson distribution of sources, which produces an angle-independent trispectrum with amplitude $t_{\rm ps}$. As shown in Paper 1, the corresponding estimator is trivial (simply a quartic product of the inverse-variance-filtered data) and will be discussed further in this work. The PolySpec lensing estimator takes a similar form to the $\tau_{\rm NL}^{\rm loc}$ estimator (involving the ‘exchange’ of a lensing field), and optimally estimates the lensing amplitude $A_{\rm lens}$ ($\propto C_{L}^{\phi\phi}$), using $T$-, $E$-, and $B$-modes. As described in Paper 1, this is analogous to the conventional quadratic lensing estimators [e.g., 86, 87, 88, 57], but naturally incorporates the $\mathrm{N}^{(1)}$ bias, first-principles ‘realization-dependent-noise’ [59], optimal polarization correlations [cf. 89], and mask- and weighting-dependent normalization.
@@
To validate the lensing estimator, we employ $150$ Gaussian realizations of temperature, polarization and lensing potential. [^20] Specifically, we use the FFP10 realizations [e.g., 90], which are publicly available on nersc. Using the lenspyx code [91], we generate non-Gaussian simulations with a known lensing amplitude $A_{\rm lens}$, which are supplemented with Gaussian noise and a Planck-mask, as before. Here, we use the hyperparameter set $\{N_{\rm disc}=100,N_{\rm fish}=50,\ell_{\rm min}=2,L_{\rm min}=2,N_{\rm side}=512\}$, fixing $L_{\rm max}=\ell_{\rm max}$ and using lensed $TT,TE,EE$ and $BB$ spectra where necessary [92], noting that the lensing estimators do not require optimization (since they do not contain radial integrals). Given that the fiducial lensing amplitude is non-zero, it is important to validate PolySpec at high resolution, thus we perform mock analyses up to $\ell_{\rm max}=2048$ (increasing $N_{\rm side}$ to $1024$ where necessary). [^21] As discussed above, it is safe to validate the other estimators at a reduced $\ell_{\rm max}$ since the errors scale with $\sigma(A)$ in the absence of a signal (assuming that the disconnected contributions are correctly subtracted).
@@ VI Summary & Next Steps
We close by elucidating our next steps. Up to this point, we have presented the inflationary templates and their optimal estimators (Paper 1) and built an efficient code to search for such signatures (this paper). In Paper 3, we will search for non-Gaussian signatures using Planck data. Motivated by the above results, we will perform both single-template and joint analyses of a range of inflationary templates, including all of the shapes considered in this work except for the poorly constrained direction-dependent forms with odd $n_{1},n_{3}$ (cf. §IV.5). We will further account for CMB lensing contamination using our $A_{\rm lens}$ template, and perform both temperature-only and joint temperature-and-polarization analyses. The result will be strong constraints on a wide variety of inflationary phenomena; these can be further improved using future CMB or large-scale structure datasets [e.g., 93, 94, 82].
@@ Acknowledgements.
We thank Giovanni Cabass, William Coulton, Adriaan Duivenvoorden, Sam Goldstein, Colin Hill and Maresuke Shiraishi for insightful discussions. We additionally thank the anonymous referee for their careful reading of the manuscript and useful comments. OHEP is a Junior Fellow of the Simons Society of Fellows, and thanks Lludwig van Beethoven for auditory assistance.OHEP would also like to thank the Center for Computational Astrophysics for their hospitality across the multiple years this set of papers took to write. The computations in this work were run at facilities supported by the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation.